If sum of the three terms of an AP is 27 and product of three terms is 504,find the terms?
Answer by mananth(16946)

let the terms be a,a+d,a+2d
a+a+d+a+2d= 27
3a+3d=27
/3
a+d=9
a(a+d)(a+2d)=504
a+d =9
a*9*(9+d)=504
a= 9-d
(9-d)(9+d)*9=504
(9+d)(9-d)=504/9
9^2-d^2=56
d^2=81-56
d^2=25
d=+/-5
a+d =9
when d = 5
a=4
the numbers are 4,9,15
The same when you take d=-5